OpenStudy (anonymous):

find the equation of the line that has a slope of 4 and passes through the point (1,4)

OpenStudy (anonymous):

find b by plugging in for x and y into the equation y=mx+b: 4=4(1)+b 0=b y=4x

OpenStudy (sgadi):

\[y=mx+c\] \[m=4\] \[y=4x+c\] \[4=4+c\Rightarrow c=0\] so equation is \[y=4x\]

OpenStudy (anonymous):

:)

OpenStudy (sgadi):

sorry ... repeated :-)

OpenStudy (anonymous):

why does b=0